OpenStudy (anonymous):

Find f'(x) for f(x)=ln(x+rad(x^(2)-1)). Can somebody please help me with this?

OpenStudy (anonymous):

\(\dagger\)..

OpenStudy (kainui):

Almost, but what you've written is kind of ambiguous in the first step of writing out the derivative. \[f'(x) \ne \frac{d}{dx}(\ln(x+\sqrt{1+x^2}))*\frac{d}{dx}(x+\sqrt{1+x^2})\] \[f'(x) = \frac{d}{dx}(\ln(x+\sqrt{1+x^2}))\] you've sort of combined the chain rule there where it shouldn't be.

OpenStudy (jhannybean):

Oh, I don't think I was een done yet. If you combine some things in my last step it could simplify even further.

OpenStudy (kainui):

So to put it explicitly, \[f(x)= h(k(x))\] \[h(x)=\ln(x) \\ k(x) = 1+\sqrt{1+x^2}\] \[f'(x)=h'(k(x))*k'(x)\] \[h'(x) = \frac{1}{x} \\ h'(k(x))=\frac{1}{1+\sqrt{1+x^2}} \\ k'(x) = \frac{x}{\sqrt{1+x^2}}\]

OpenStudy (jhannybean):

Omg, i also realized I mixed up a lot of my signs :|

OpenStudy (jhannybean):

\[\color{red}{\sf f(x) = \ln(x+\sqrt{x^2-1})}\]Just pretend the derivative i am taking to the right is actually a fraction ontop of the derivative of the log function.\[\sf f'(x) = \frac{d}{dx}(\ln(x+\sqrt{x^2-1}) \cdot \frac{d}{dx} (x+\sqrt{x^2-1})\]\[\sf f'(x) = \frac{1}{x+\sqrt{x^2-1}}\cdot \left(1+\frac{1}{2}(x^2-1)^{-1/2}\cdot 2x\right)\]\[\sf f'(x)= \left[\frac{1}{x+\sqrt{x^2-1}}\right]\cdot \left(1+\frac{x}{(x^2-1)^{1/2}}\right)\]\[\sf f'(x)=\left[\frac{\left(1+\frac{x}{\sqrt{x^2-1}}\right)}{\sqrt{x^2-1} \left(\frac{x}{\sqrt{x^2-1}}+1\right)}\right]\]Realize, once again, that i just took the numerator and separated it from the denominator at first to simplify it for myself, but I just put it back ontop.If we follow the format \(\frac{a}{b(a)} =\frac{1}{b}\) we will end up with :\[\sf \boxed{\large f'(x) = \frac{1}{\sqrt{x^2-1}}}\] :)

OpenStudy (jhannybean):

\[\sf f'(x)= \left[\frac{1}{x+\sqrt{x^2-1}}\right]\cdot \left(1+\frac{x}{\sqrt{x^2-1}}\right)\\ =\left[\frac{1}{x+\sqrt{x^2-1}}\right]\cdot \left(\frac{\sqrt{x^2-1}+x}{\sqrt{x^2-1}}\right)\\ =\frac{1}{\sqrt{x^2-1}}\] Another sorter method by @UnkleRhaukus :)

OpenStudy (kainui):

\(\color{blue}{\text{Originally Posted by}}\) @Jhannybean \[\color{red}{\sf f(x) = \ln(x+\sqrt{x^2-1})}\]Just pretend the derivative i am taking to the right is actually a fraction ontop of the derivative of the log function.\[\sf f'(x) = \frac{d}{dx}(\ln(x+\sqrt{x^2-1}) \cdot \frac{d}{dx} (x+\sqrt{x^2-1})\] \(\color{blue}{\text{End of Quote}}\) Don't write it if it's wrong. That's wrong. Calculus is confusing enough.

OpenStudy (jhannybean):

Asked a number of my math professors about the format on this problem, and was told there was nothing wrong with writing it that way.

OpenStudy (mendicant_bias):

Some messages look like they are deleted/missing for me. Kainui, was your reply to a post that the OP deleted?

OpenStudy (mendicant_bias):

PS is this mathjax or what even is this? This doesn't look like LaTeX in its typical format. Is this just upright math LaTeX 24/7 or a different font or something?

OpenStudy (jhannybean):

It's just a different font. `/sf`

OpenStudy (jhannybean):

Hmm.. maybe it would have been better if I let \(u=x+\sqrt{x^2-1}\) and \(du=1+\frac{x}{\sqrt{x^2-1}}\) Or something. Oh freaking well.

OpenStudy (jhannybean):

It's really bothering me now.

OpenStudy (mendicant_bias):

I know it's somewhere in my Calc book, there's a better way to do what you're trying to illustrate using composite functions, but I haven't looked it up yet; might eventually.

OpenStudy (jhannybean):

You mean conjugates?

OpenStudy (anonymous):

Hmm...

OpenStudy (anonymous):

I think what you SHOULD have written was: \[ f'(x)=\frac{d}{d\left(x+\sqrt{1+x^2}\right)}\ln\left(x+\sqrt{1+x^2}\right)\cdot \frac{d}{dx}\left(x+\sqrt{1+x^2}\right) \]Or if you had said \(u = \left(x+\sqrt{1+x^2}\right)\) then: \[ f'(x)=\frac{d}{du}\ln\left(u\right)\cdot \frac{d}{dx}\left(x+\sqrt{1+x^2}\right) \]

OpenStudy (anonymous):

This ultimately would mean: \[ \frac{df}{dx}=\frac{df}{du}\frac{du}{dx} \]However, it's a minor notation error. What you were trying to say was correct, but what you actually wrote was a bit incorrect.

OpenStudy (anonymous):

Technically: \[\begin{split} f'(x)&=\frac{d}{dx}\ln\left(x+\sqrt{1+x^2}\right)\cdot \frac{d}{dx}\left(x+\sqrt{1+x^2}\right) \\ &= f'(x)\frac{d}{dx}\left(x+\sqrt{1+x^2}\right) \end{split} \]Which doesn't make sense.

OpenStudy (jhannybean):

This makes more sense. Thank you.